OpenStudy (anonymous):

Write the equation of the line that is perpendicular to the line 3x + y = 7 and passes through the point (6, −1).

OpenStudy (anonymous):

First find the gradient of the line: 3x + y = 7 y= -3x + 7 Gradient is given by the coefficient of x which is -3 in this case. The gradient perpendicular to -3 is 1/3 Given the points (6,-1) and gradient,m =1/3 Substitute it into the equation for a line y=mx+c -1 = 1/3(6) +c c= -3 Eqn of line, y= 1/3 x-3